Return on Equity

Commonwealth Construction (CC) needs $3 million of assets to get started, and it expects to have a basic earning power ratio of 35%. CC will own no securities, so all of its income will be operating income. If it so chooses, CC can finance up to 60% of its assets with debt, which will have an 10% interest rate. If it chooses to use debt, the firm will finance using only debt and common equity, so no preferred stock will be used. Assuming a 30% tax rate on all taxable income, what is the difference between CC's expected ROE if it finances these assets with 60% debt versus its expected ROE if it finances these assets entirely with common stock? Round your answer to two decimal places.
Answer #1

Solution:

Given Total Assets = $3,000,000, Interest rate = 10%, tax rate = 30%, Basic earning power ratio = 35% = EBIT/Total assets, so

EBIT = 0.35 ($3,000,000) = $1,050,000, D/A = 0.6, So equity = $1,200,000

D/A = 0% D/A = 60%

EBIT $1,050,000 $1,050,000

Interest 0 $180,000 ( $3,000,000*0.6*10%)

EBT $1,050,000 $870,000

Tax (30%) $315,000 $261,000

Net Income $735,000 $609,000

ROE = Net Income/Equity

ROE = 735,000/3,000,000 = 24.5% ROE = 609,000/1,200,000 = 50.75%

Difference in ROE = 50.75% - 24.5% = 26.25%
Answer #2
27.6% answer is common equity
Answer #3
64.4% answer is debt equity